Hi there,
My tenant owes me Rs. 6,199,267 in rental arrears which started in January 2012 (2 and a 1/2 years). He is leaving on June 30th and has agreed to pay me, the total in full, at an interest rate of 11%. What would be the total outstanding? How would it be compounded? monthly, daily, or yearly? If possible, could you show me all three and how you did it? Thank you so Much for your help.
GOD BLESS YOU! :-)
Dinuke

Answer by ankor@dixie-net.com(22740)

My tenant owes me Rs. 6,199,267 in rental arrears which started in January 2012 (2 and a 1/2 years). He is leaving on June 30th and has agreed to pay me, the total in full, at an interest rate of 11%. What would be the total outstanding? How would it be compounded?
:
Make it simple
:
Divide the total amt by 2; 6199267/2 = 3099633.5
We can assume that half the time he owed less than this amt and half the time he owed more. Take this as the average amt he owed for 2.5 yrs
Find the interest
I = 3099633.5 * .11 * 2.5 = 852399 in interest
add to amt owed
6199267 + 852399 = $7,051.666 total owed